Return  Xy-pic  Young  Myhelp2

Make your own HELP!

1. Copy the source file ./HyperTeX/input7.tex of this DVI file

as a name input1.tex.

2. Copy myhyper.sty and keyin.sty in the same directory.

3. Rewrite input1.tex as you like after the line

    %%% Rewrite from Here %%%

4. For example, the result of the source text

  \Q{\alpha}
  \E{\yen}
  \keyin{\widehat{}}{$\widehat{x+y}$}
  \keyin{\begin{quote} 0d 0a   0d 0a\end{quote}}{quote}

in input1.tex is as follows

α ¥ { \atop x + y} quote

  Here \keyin{#1}{#2} means

#2 is a hot spot

#1 is a string to be copied to clipboard and 0d means the character with code 0d in Hex and is translated into

  \E{#1} means \keyin{#1}{#1}.

If #1 contains multiple lines, it should be \K{#1\K}

  \Q{#1} means \keyin{#1}{$#1$}.

See ./HyperTex/input.tex for more examples.

5. Make input1.dvi and copy it to ./HyperTeX/.

6. Then by clicking MyHelp in Help TeX, input1.dvi is shown.

Remark. In the above, input1.tex may be input0.tex, …,input6.tex. If foo?.dvi is indicated by HyperTEX, then ? is replaced by 0,1,…,9 and the corresponding DVI files are searched in this order. Then the DVI file first found is displayed.

Hence input1.dvi has priority over input7.dvi and input1.dvi will not be changed even if a new package of dviout is overwritten.

Top